A 982.0 N force is exerted on a circular surface with a radius of 2.0 cm. What is the resulting pressure on the surface? Express your answer in units of psi (pounds per square inch), a common unit of pressure.
Hint: In order to make the unit conversions, you will have to use the fact that:
1 N = 0.2248 lb (pounds) and
1 in = 2.540 cm
dont know y iam not getting the right answer plz help.

Answer by rothauserc(4718)

Pressure(P) = Force(F)/Area(A)
:
In the foot-pound-second (FPS) system, the P units are pounds per square inch, or psi.
:
F in pounds = 982.0 * 0.2248 = 220.7536
:
radius in inches = 2/2.540 = 0.7874
:
A in square inches of the circular surface = pi * (0.7874)^2 = 1.9478
:
*************************************
P in psi = 220.7536/1.9478 = 113.3348
*************************************
: